OpenStudy (anonymous):

A recent poll found that 45% of eligible voters are planning to vote in favor of a new by-law. Suppose you randomly survey six voters. What is the probability that at least three of the voters plan to vote in favor of the new by-law? 13.2% 25.5% 30.3% 55.8%

OpenStudy (anonymous):

@amistre64 this one is for you

OpenStudy (anonymous):

I think this is binomial distribution so you have to use the formula \[\sum_{k=0}^{X}\left(\begin{matrix}n \\ k\end{matrix}\right)p^k(1-p)^(n-k)\] n = number of trials p = probability of success k = number of successes

OpenStudy (anonymous):

Or you can use the binomial distribution tables to find \[P(X \ge 3)\] which is also \[1 - P(X \le 2)\]

OpenStudy (amistre64):

i was thinking the P(X>=3) method :)